Draw a lewis structure for H2S include lone pair electrons


1
Expert's answer
2021-01-06T05:55:07-0500

Q153168

Draw a lewis structure for H2S include lone pair electrons


Solution:


Step 1: Find the total valence electrons in H2S.

The group number represents the number of valence electrons of the elements present in that group. Hydrogen is present in the 1 A group and S is present in the 6 A group.

So '# of valence electrons in H' is 1 and '# of valence electrons in S' is 6.


There are two Hydrogen atoms and one Sulphur atom in each H2S molecule,

So total valence electrons = 2 (1) + 6 = 8 .



Step 2: Draw the skeletal structure of H2S.


Lewis structure of a compound is a structural representation that represents the valence electrons of atoms within a molecule. 

The given compound contains atoms of only H and S. H can form only a single bond.

So we will keep S in the middle with both H attached to it by a single bond,




Each bond requires two electrons. So total electrons used in showing two bonds is 4.


# of valence electrons left with us after showing bonds = 8 - 4 = 4.


Step 3: Complete the octet of the S atom.


Hydrogen needs two electrons for completing its outermost shell. Sulfur needs 8 electrons for completing its outermost shell.

Each bond represents two electrons, so the outermost shell of hydrogen is already completed.

We will show the remaining 4 electrons on the S atom.




There are two bonding and two non-bonding pairs present in this molecule.



This is the Lewis structure of H2S molecule.

The molecular geometry of H2S will be bent or angular.